How long will it take $200 to double if it earns 7%? Compounding occurs once a year? Time Value of Money. I calculated the Present Value by taking the # of 1 yrs using my calculator. 1N, 7 I/Y, 200, FV, CPT, PV=$186.92 for the 7% rate. I am stuck after that.
Answer by Alan3354(69443)

400 = 200*(1 + 0.07)^t t in years
1.07^t = 2
t*log(1.07) = log(2)
t = log(2)/log(1.07)
t =~ 10.244 years
--> 11 years
----------------
The amount doesn't matter, it's doubled.
1.07^11 = 2.10485